# How important is the dielectric constant in water modeling? Evaluation of the performance of the TIP4P/ force field and its compatibility with the Joung-Cheetham NaCl model

### v1

@@ 3.1 Dielectric constant
3.2 Melting point
The melting point of ice Ih was estimated using the direct coexistence method 51 where two phases at coexistence are brought together across an interface. Based on the time evolution of number of liquid molecules, identified by the CHILL+ algorithm 52, we have estimated the melting/freezing rates. Afterward, these rates were plotted as a function of temperature and the melting point was located by interpolation of results to zero rate. The simulations performed in the ensemble lasted up to ns to unambiguously determine the melting/freezing modes.
3.3 Infrared spectrum
The infrared spectra of rigid point charge models are estimated following a procedure suggested by Skinner and collaborators 53, 54, 55, 56. In this method the vibrational frequencies of water clusters calculated from ab-initio calculations are correlated to the electric fields generated by a rigid point charge model. In this way a spectroscopic map is created which allows one to predict the infrared spectra from the electric field generated by a force field 57. Here, we used the time averaging approximation, whereby the instantaneous vibrational frequencies as obtained from the spectroscopic map are averaged over a period of the order of the librational timescale. The spectrum is then predicted as the distribution of such frequencies over the full simulation, according to 54:
where is the time averaged frequencies. Here we focus on the OD tension of a solution of deuterated water (HDO) in H2O, under the assumption that any one single molecule of the system is representative of the electric field felt by the solvated HDO molecule 58. This choice is convenient, because the OD chromophore does not couple to the neighboring OH chromophores, due to the large separation of the frequencies. As a result, we need not embark into the complications of inter or intra molecular coupling 54, 55, 56. In order to estimate instantaneous frequencies to each molecular environment as sampled from the simulations, we use the OD stretch map parametrized for the TIP4P model 56. A number of different studies point to the robustness of this mapping to change of model potentials and local environmental conditions 56, 57, 59, 58 Instantaneous frequencies are time averaged over 100 ps, as recommended in Ref.60 Averages are collected from NVE simulations with 256 molecules, using a time step of 0.5 fs in order to guarantee an accurate calculation of the high frequency dynamics 61.
3.4 Transport properties
@@
where is the position of a water molecule at time , and the triangular brackets denote a thermal average over all time origins. The diffusion coefficient is then related to the slope of the MSD as . To account for the finite-size effects we have included the Yeh-Hummer (YH) correction 62, defined as
@@
where , with represent the components of the stress tensor. To improve the statistics, we exploit both the off-diagonal and the diagonal components of the stress tensor. This is allowed provided the off-diagonal elements are weighted by the adequate factors 63. Taking this into account, the shear viscosity is calculated as:
@@
The simulation scheme for both transport properties involved two steps. In the first one, auxiliary simulations were performed for ns, allowing to estimate the average box size and density of the fluid. Afterwards, the system was rescaled to the average dimensions estimated before and production runs lasting up to ns in the ensemble were launched. This avoids possible spurious effects of the barostat in the dynamics of the system.
3.5 Properties of ice polymorphs
The density of ice II, V and VI at selected conditions was calculated using a Monte Carlo code used previously in the determination of the phase diagram of ice and the parametrization of TIP4P/2005 and TIP4P/Ice models13, 64, 15. Details of the code may be found in Ref.65. Details on the preparation of initial configurations of the ice polymorphs are described in Ref.13, 66 The dielectric constant of ice Ih was calculated using a ring rotation algorithm as described in Ref.67, 68, 69
3.6 Molecular dynamics setup
| System | Number of molecules | |
@@ 3.6 Molecular dynamics setup
The molecular dynamics simulations are performed using the LAMMPS package 70. Water was modeled using the TIP4P/ force field 39 and for selected properties we have also performed calculations for the TIP4P/2005 model. The number of molecules and box sizes varied in different simulation sets and are summarized in Table 1. Trajectories are evolved using the velocity-Verlet integrator using a time step of fs. For pure water, the equations of motion were solved with the SHAKE algorithm, with the dummy atom defined implicitly according to the the pair_style lj/cut/tip4p/cut command. Unfortunately this command does not allow for the correct calculation of electrostatic energy in the presence of ions. Therefore, in the case of solutions a slightly less efficient quaternion-based rigid body dynamics was used 71. Both the temperature and pressure were maintained using a three chains Nosé-Hoover algorithm with a damping factor ps. Tail corrections were included. Dispersion interactions were truncated at the separation distance of Å. Long-range electrostatics were computed using the particle–particle particle-mesh method 72. The charge structure factors were evaluated with the fourth-order interpolation scheme and a grid spacing of 1 Å.
3.7 Force field parameters
In this work, we have performed simulations using the TIP4P/ water model 39 which has been compared to the two other non-polarizable force fields - TIP4P/2005 64 and OPC 38. The entire set of parameters, required for launching the simulations is summarized in Table 2.
| Model | Charge (e) | (Å) | (kcal/mol) | (Å) | (Å) | (D) | ||
@@ 3.7 Force field parameters
4 Results and Discussion
4.1 Dielectric constant
Let us begin with the discussion related to the static dielectric constant of the TIP4P/ water model. In the initial paper, Fuentes-Azcatl et al. 39 argued that Lennard-Jones and Coulombic interactions behave independently. This allows to choose the charge distribution to reproduce the dielectric constant and then determine the LJ parameters to reproduce the temperature of maximum density (TMD) and other properties. In fact, they have shown that this approach was successful and the developed water model could reproduce experimental values of both properties.
@@ 4.1 Dielectric constant
At first thought, one could suspect that such a large increase of the dielectric constant can be achieved by artificially increasing the molecular dipole with too large point charges. This would then seriously compromise the accuracy of the electric field generated by the molecules, thus spoiling the actual force field. Accordingly, the good agreement of dielectric constants would be at the cost of seriously deteriorating prediction of the remaining properties. Actually, inspection of the model parameters (c.f. Table 2) shows that the TIP4P/ has point charges that are just 5% smaller than those of the TIP4P/2005. The crucial difference between the models appears to be the placement of the dummy particle, which is located much closer to the oxygen site in TIP4P/. The overall effect of this change is to increase the single molecule dipole moment by merely 5%. Accordingly, the large increase in the dielectric constant cannot be attributed to a crude scaling of the dipole.
@@
A very stringent test for non-polarizable models is to check the drop of the dielectric constant upong freezing water. In experiments, the dielectric constant at the melting point of ice increases from 78 for ice cold water, to 107 for ice Ih. This large increase has never been predicted for any of the point charge models. In fact, these models, on the contrary, predict a significant drop of the dielectric constant on freezing 75, 69. TIP4P/ is no exception. Our calculations show that for ice at 273 K, the dielectric constant decreases down to 43, due to a very large reduction of the Kirkwood factor, which drops to . On the other hand, TIP4P/2005 exhibits a dielectric constant of , and now achieves a Kirkwood factor larger than that of TIP4P/. Whence, although the two models exhibit similar dielectric constants at this temperature (where ice Ih is actually a metastable phase for both models), we see that the behavior of TIP4P/2005 is qualitatively somewhat better than that of TIP4P/, as it predicts a much smaller drop of the dielectric constant upon freezing. In fact, the behavior of TIP4P/ somewhat resembles that of the SPC/E model, where the Kirkwood factor decreases from 3.68 for water at ambient temperature to barely 1.74 for ice 75. This similarity is reasonable, since TIP4P/ has the dummy site displaced from the oxygen by a much smaller amount than TIP4P/2005, and is therefore more closely related to SPC/E. If this analogy holds, TIP4P/ is not a reliable model for the study of ice polymorphs. This statement will be assessed later on.
4.2 Infra Red spectrum
@@
Since TIP4P/2005 and TIP4P/ seem to produce rather different charge distributions, it is instructive to move on to study high frequency dipolar fluctuations of condensed water, which are very sensitive to the local electric fields generated on the water atoms 53, 76, 54. The OH stretch band of the Infrared spectrum is a particularly convenient probe. However, in order to avoid complications due to inter and intra molecular couplings, here we consider the OD stretch of an isolated HDO molecule dissolved in H2O, which behaves as an uncoupled chromophore 53, 76, 60, 61.
@@
Of course, rigid models do not exhibit an explicit OD stretch at all, but it is possible to produce a synthetic spectrum from the study of the local electric fields generated on the Deuterated atoms 53, 54, 55, 56. Although this is an empirical approach, which maps ab-initio results for the OD stretch to the electric field of a reference TIP4P model, it has been argued that the mapping is transferable and can be exploited to assess the performance of different force fields 57, 59, 58.
@@
The results of the OD spectrum are shown in Figure 2. From panel a) we see that TIP4P/ does a much better job than TIP4P/2005 at describing the low frequency side of the OD band, but has a significantly larger band width, and does therefore not reproduce the high frequency side of the band accurately. On the other hand, the spectrum predicted by TIP4P/2005 has a band center that is some decades of cm-1 blue shifted with respect to experiment, but shows a band width closer to the experimental spectrum.
@@
Both models do a good job at reproducing the blue shift of the OD stretch as temperature increases, with TIP4P/ yielding a band center in closer agreement with experiment. However, the slope of the band center as a function of temperature is given more accurately in the TIP4P/2005 model.
@@ 4.3 Densities
4.4 Melting point
A well known limitation of point charge models is the difficulty to predict simultaneously both the melting point and the temperature of maximum density. TIP4P/2005, which is otherwise quite good as a model of liquid water, predicts the melting point at about 250 K. TIP4P/ is no exception, and previous reports of its melting point yield K. However, being aware of the uncertainties in the evaluation of the due to the finite-size effects 78, we have repeated the estimation of using the direct coexistence method with a large system size comprising of water molecules in total. The secondary prism face (pII) of ice was exposed to water as it exhibits the fastest kinetics 79 and the evolution of number of liquid molecules as a function of time was monitored for several different temperatures. This can be seen in Figure 4-a.
We can estimate the freezing/melting rates from the slope of the number of liquid molecules as a function of time. These are then plotted as a function of temperature, and the melting point is estimated as the interpolation of this curve to zero melting rate 65. The freezing/melting rates are plotted in Figure 4-b. The intersection of these rates with the x-axis provides an estimated melting temperature of K. This value is smaller than the recently estimated melting point of the TIP4P/2005 model of K 78. For point charge models with similar dipole moment, the location of the melting point correlates with the quadrupole moment, 80. As shown in Table 2, TIP4P/ has lower value of than the TIP4P/2005, explaining the shift in melting point by about K.
@@ 4.5 Transport properties
By virtue of the Stokes-Einstein relation, it is expected that the shear viscosity values should also be equivalently well reproduced by the TIP4P/ water model as it is inherently coupled to the self-diffusion coefficient. Indeed, it is the case as can be seen in Figure 6 where the results overlap with the data from Montero de Hijes et al. 81 for the TIP4P/2005 model. In Figure 6-b a magnified region is presented.
@@
In Figure 6-c, the relation of shear viscosity with pressure at three different temperatures is shown. Remarkably, the shear viscosities at K are around 10% too small as compared to the experimental data, however, the trend in which the viscosity reaches its minimum and increase again with pressure is preserved. Nevertheless, this is also the case for the TIP4P/2005 water model as demonstrated in Figure 2 in Ref. 81. As the temperature increases, the agreement with experimental results improves.
4.6 Comparison of the models
@@
So far, for selected properties we have seen that the TIP4P/ water model is essentially identical to the TIP4P/2005 force field in a wide range of properties, with the exception that the former also reproduces experimental values of the dielectric constant. Therefore, we are now in position to perform a similar comparison as has been recently done by Sedano et al. (TIP4P/2005 vs OPC models), using the VA-test method 16.
@@
The results for different properties are shown in Table 3. The comparison is done exactly as in Ref. 41, except for the heat capacities at constant pressure, which were not calculated in this work. All the results for the TIP4P/2005 and OPC water models are taken from that reference, whereas properties of the TIP4P/ are calculated by ourselves, unless marked as an asterisk. In the latter case, the value for a given property is taken from Ref. 39. The tolerances allowed for variations of properties in the rating scheme are as those reported in Table 2 of Ref. 16.
@@
Despite the wide range of properties tested in the VA benchmark, both models perform quite well. TIP4P/2005 is superior for the prediction of the melting and critical points, but is very similar to TIP4P/ for all other properties, including ice polymorph densities, liquid densities, shear viscosities and diffusion coefficients. The overall performance is almost the same for both models (7.56 for TIP4P/2005 vs 7.54 for TIP4P/) while the OPC is clearly inferior (with a score of 6.26). As a bonus, TIP4P/ predicts accurately the dielectric constants, compared to TIP4P/2005 which provides results that are ca 25 % too small. Surprisingly, this large increase of is achieved with only a minor increase of the molecular dipole moment. The implication is that orientation correlations of the TIP4P/ model must be very different from those predicted by TIP4P/2005, yet the prediction of densities and transport properties in both models remain very similar. It would seem that changes in the Lennard-Jones parameters from one model to the other compensate somehow the differences in charge distributions. The question then remains, whether one of the models is predicting the actual electric field about a water molecule better than the other, but this cannot be resolved from our results based on the properties of pure water.
@@
| Enthalpy of phase change (kcal/mol) | ||||
| 1.44 | 1.13 | 0.971 | 1.07 | |
| 10.52 | 11.99 | 12.9 | ||
@@
5 Electrolyte properties of TIP4P/ model
In order to elucidate which of the two models is more consistent, it is required to test their performance beyond pure water properties. Particularly, a strategy to test which of the two charge distributions is electrostatically more significant is to test the models under a strong electric fields, such as that generated locally in electrolyte solutions. A priori, one can tell that TIP4P/ could perform better for the solubilities of salts, as it predicts accurately the dielectric constant. However, if this agreement is at the cost of spoiling the force field, the remaining solution properties will not be predicted correctly.
5.1 Joung-Cheetham NaCl model
In the present work, we have decided to use the NaCl parameters of the force field that has been developed by Joung and Cheetham (JC) 30. The parameters used in conjunction with the TIP4P/ water model are shown in Table 4.
The motivation behind the choice of the JC model is three-fold. Firstly, the model has been parametrized to accurately reproduce the properties of the NaCl crystal and its melt. This constrains the parameters and avoids possible inconsistencies that could result from fitting to the solution properties alone. Secondly, as a corollary of the former comment, the C model sets unit charges to the Na+ and Cl- ions, which is the text book expectation for charge distributions of dissolved ions. Thirdly, the model parameters have not been optimized for either TIP4P/2005 or TIP4P/. This is convenient, as we can therefore check the performance of the water models without any a priori bias. The comparison will therefore allow us to check whether (i) the model eletrolyte is transferable to other water models and (ii) reproducing the experimental values of dielectric constant improves the results.

### v2

@@ III.1 Dielectric constant
III.2 Melting point
The melting point of ice Ih was estimated using the direct coexistence method Garc´ıa Fernández, Abascal, and Vega (2006) where two phases at coexistence are brought together across an interface. Based on the time evolution of a number of liquid molecules, identified by the CHILL+ algorithm Nguyen and Molinero (2015), we have estimated the melting/freezing rates. Afterwards, these rates were plotted as a function of temperature and the melting point was located by interpolation of results to zero rate. The simulations performed in the ensemble lasted up to ns to unambiguously determine the melting/freezing modes.
III.3 Infrared spectrum
The infrared spectra of rigid point charge models are estimated following a procedure suggested by Skinner and collaborators Corcelli, Lawrence, and Skinner (2004); Auer and Skinner (2007, 2008); Gruenbaum et al. (2013). In this method, the vibrational frequencies of water clusters calculated from ab initio calculations are correlated to the electric fields generated by a rigid point charge model. In this way a spectroscopic map is created which allows one to predict the infrared spectra from the electric field generated by a force field Takayama, Otosu, and Yamaguchi (2023). Here, we used the time averaging approximation, whereby the instantaneous vibrational frequencies as obtained from the spectroscopic map are averaged over a period of the order of the librational timescale. The spectrum is then predicted as the distribution of such frequencies over the full simulation, according to Auer and Skinner (2007):
where is the time averaged frequencies. Here we focus on the OD tension of a solution of deuterated water (HDO) in H2O, under the assumption that any one single molecule of the system is representative of the electric field felt by the solvated HDO molecule Takayama, Otosu, and Yamaguchi (2024). This choice is convenient, because the OD chromophore does not couple to the neighboring OH chromophores, due to the large separation of the frequencies. As a result, we need not embark on the complications of inter or intra molecular coupling Auer and Skinner (2007, 2008); Gruenbaum et al. (2013). In order to estimate instantaneous frequencies to each molecular environment as sampled from the simulations, we use the OD stretch map parametrized for the TIP4P model Gruenbaum et al. (2013). A number of different studies point to the robustness of this mapping to change of model potentials and local environmental conditions Gruenbaum et al. (2013); Takayama, Otosu, and Yamaguchi (2023); Yamaguchi, Takayama, and Otosu (2023); Takayama, Otosu, and Yamaguchi (2024) Instantaneous frequencies are time averaged over 100 ps, as recommended in Ref.Kananenka and Skinner (2018) Averages are collected from NVE simulations with 256 molecules, using a time step of 0.5 fs in order to guarantee an accurate calculation of the high frequency dynamics Dicu-Gohoreanu (2024).
III.4 Transport properties
@@
where is the position of a water molecule at time , and the triangular brackets denote a thermal average over all time origins. The diffusion coefficient is then related to the slope of the MSD as . To account for the finite-size effects we have included the Yeh-Hummer (YH) correction Yeh and Hummer (2004), defined as
@@
where , with represent the components of the stress tensor. To improve the statistics, we exploit both the off-diagonal and the diagonal components of the stress tensor. This is allowed provided the off-diagonal elements are weighted by the adequate factors Daivis and Evans (1994). Taking this into account, the shear viscosity is calculated as:
@@
The simulation scheme for both transport properties involved two steps. In the first one, auxiliary simulations were performed for ns, allowing us to estimate the average box size and density of the fluid. Afterwards, the system was rescaled to the average dimensions estimated before and production runs lasting up to ns in the ensemble were launched. This avoids possible spurious effects of the barostat in the dynamics of the system.
III.5 Properties of ice polymorphs
The density of ice II, V and VI at selected conditions was calculated using a Monte Carlo code used previously in the determination of the phase diagram of ice and the parametrization of TIP4P/2005 and TIP4P/Ice modelsSanz et al. (2004); Abascal and Vega (2005); Abascal et al. (2005). Details of the code may be found in Ref.Lukasz Baran and MacDowell (2024). Details on the preparation of initial configurations of the ice polymorphs are described in Ref.Sanz et al. (2004); MacDowell et al. (2004) The dielectric constant of ice Ih was calculated using a ring rotation algorithm as described in Ref.Rick and Haymet (2003); Rick (2005); MacDowell and Vega (2010)
III.6 Molecular dynamics setup
| System | Number of molecules | (Å3) |
@@ III.6 Molecular dynamics setup
The molecular dynamics simulations are performed using the LAMMPS package Thompson et al. (2022). Water was modeled using the TIP4P/ force field Fuentes-Azcatl and Alejandre (2014) and for selected properties we have also performed calculations for the TIP4P/2005 model. The number of molecules and box sizes varied in different simulation sets and are summarized in Table 1. Trajectories are evolved using the velocity-Verlet integrator using a time step of fs. For pure water, the equations of motion were solved with the SHAKE algorithm, with the dummy atom defined implicitly according to the pair_style lj/cut/tip4p/cut command. Unfortunately, this command does not allow for the correct calculation of electrostatic energy in the presence of ions. Therefore, in the case of solutions, a slightly less efficient quaternion-based rigid body dynamics was used Kamberaj, Low, and Neal (2005). Both the temperature and pressure were maintained using a three chains Nosé-Hoover algorithm with a damping factor ps. Tail corrections were included. Dispersion interactions were truncated at the separation distance of Å. Long-range electrostatics were computed using the particle–particle particle-mesh method Darden, York, and Pedersen (1993). The charge structure factors were evaluated with the fourth-order interpolation scheme and a grid spacing of 1 Å.
III.7 Force field parameters
In this work, we have performed simulations using the TIP4P/ water model Fuentes-Azcatl and Alejandre (2014) which has been compared to the two other non-polarizable force fields - TIP4P/2005 Abascal and Vega (2005) and OPC Izadi, Anandakrishnan, and Onufriev (2014). The entire set of parameters, required for launching the simulations is summarized in Table 2.
| Model | Charge (e) | (Å) | (kJ/mol) | (Å) | (Å) | (D) | ||
@@ III.7 Force field parameters
IV Results and Discussion
IV.1 Dielectric constant
Let us begin with the discussion related to the static dielectric constant of the TIP4P/ water model. In the initial paper, Fuentes-Azcatl et al. Fuentes-Azcatl and Alejandre (2014) argued that Lennard-Jones and Coulombic interactions behave independently. This allows us to choose the charge distribution to reproduce the dielectric constant and then determine the LJ parameters to reproduce the temperature of maximum density (TMD) and other properties. In fact, they have shown that this approach was successful and the developed water model could reproduce experimental values of both properties.
@@ IV.1 Dielectric constant
At first thought, one could suspect that such a large increase of the dielectric constant can be achieved by artificially increasing the molecular dipole with too large point charges. This would then seriously compromise the accuracy of the electric field generated by the molecules, thus spoiling the actual force field. Accordingly, the good agreement of dielectric constants would be at the cost of seriously deteriorating prediction of the remaining properties. Actually, inspection of the model parameters (cf. Table 2) shows that the TIP4P/ has point charges that are just 5% smaller than those of the TIP4P/2005. The crucial difference between the models appears to be the placement of the dummy particle, which is located much closer to the oxygen site in TIP4P/. The overall effect of this change is to increase the single molecule dipole moment by merely 5%. Accordingly, the large increase in the dielectric constant cannot be attributed to a crude scaling of the dipole.
@@
A very stringent test for non-polarizable models is to check the drop of the dielectric constant upon freezing water. In experiments, the dielectric constant at the melting point of ice increases from 78 for ice cold water, to 107 for ice Ih. This large increase has never been predicted for any of the point charge models. In fact, these models, on the contrary, predict a significant drop of the dielectric constant on freezing Aragones, MacDowell, and Vega (2011b); MacDowell and Vega (2010). TIP4P/ is no exception. Our calculations show that for ice at 273 K, the dielectric constant decreases down to 43, due to a very large reduction of the Kirkwood factor, which drops to . On the other hand, TIP4P/2005 exhibits a dielectric constant of , and now achieves a Kirkwood factor larger than that of TIP4P/. Whence, although the two models exhibit similar dielectric constants at this temperature (where ice Ih is actually a metastable phase for both models), we see that the behavior of TIP4P/2005 is qualitatively somewhat better than that of TIP4P/, as it predicts a much smaller drop of the dielectric constant upon freezing. In fact, the behavior of TIP4P/ somewhat resembles that of the SPC/E model, where the Kirkwood factor decreases from 3.68 for water at ambient temperature to barely 1.74 for ice Aragones, MacDowell, and Vega (2011b). This similarity is reasonable, since TIP4P/ has the dummy site displaced from the oxygen by a much smaller amount than TIP4P/2005, and is, therefore, more closely related to SPC/E. If this analogy holds, TIP4P/ is not a reliable model for the study of ice polymorphs. This statement will be assessed later on.
IV.2 Infra Red spectrum
@@
Since TIP4P/2005 and TIP4P/ seem to produce rather different charge distributions, it is instructive to move on to study high frequency dipolar fluctuations of condensed water, which are very sensitive to the local electric fields generated on the water atoms Corcelli, Lawrence, and Skinner (2004); Corcelli and Skinner (2005); Auer and Skinner (2007). The OH stretch band of the Infrared spectrum is a particularly convenient probe. However, in order to avoid complications due to inter and intra molecular couplings, here we consider the OD stretch of an isolated HDO molecule dissolved in H2O, which behaves as an uncoupled chromophore Corcelli, Lawrence, and Skinner (2004); Corcelli and Skinner (2005); Kananenka and Skinner (2018); Dicu-Gohoreanu (2024).
@@
Of course, rigid models do not exhibit an explicit OD stretch at all, but it is possible to produce a synthetic spectrum from the study of the local electric fields generated on the Deuterated atoms Corcelli, Lawrence, and Skinner (2004); Auer and Skinner (2007, 2008); Gruenbaum et al. (2013). Although this is an empirical approach, which maps ab initio results for the OD stretch to the electric field of a reference TIP4P model, it has been argued that the mapping is transferable and can be exploited to assess the performance of different force fields Takayama, Otosu, and Yamaguchi (2023); Yamaguchi, Takayama, and Otosu (2023); Takayama, Otosu, and Yamaguchi (2024).
@@
The results of the OD spectrum are shown in Figure 2. From panel a) we see that TIP4P/ does a much better job than TIP4P/2005 at describing the low frequency side of the OD band, but has a significantly larger band width, and does therefore not reproduce the high frequency side of the band accurately. On the other hand, the spectrum predicted by TIP4P/2005 has a band center that is some decades of cm-1 blue shifted with respect to experiment but shows a band width closer to the experimental spectrum.
@@
Both models do a good job at reproducing the blue shift of the OD stretch as temperature increases, with TIP4P/ yielding a band center in closer agreement with the experiment. However, the slope of the band center as a function of temperature is given more accurately in the TIP4P/2005 model.
@@ IV.3 Densities
IV.4 Melting point
A well-known limitation of point charge models is the difficulty to predict simultaneously both the melting point and the temperature of maximum density. TIP4P/2005, which is otherwise quite good as a model of liquid water, predicts the melting point at about 250 K. TIP4P/ is no exception, and previous reports of its melting point yield K. However, being aware of the uncertainties in the evaluation of the due to the finite-size effects Conde, Rovere, and Gallo (2017), we have repeated the estimation of using the direct coexistence method with a large system size comprising of water molecules in total. The secondary prism face (pII) of ice was exposed to water as it exhibits the fastest kinetics Nada and Furukawa (2005) and the evolution of a number of liquid molecules as a function of time was monitored for several different temperatures. This can be seen in Figure 4-a.
We can estimate the freezing/melting rates from the slope of the number of liquid molecules as a function of time. These are then plotted as a function of temperature, and the melting point is estimated as the interpolation of this curve to zero melting rate Lukasz Baran and MacDowell (2024). The freezing/melting rates are plotted in Figure 4-b. The intersection of these rates with the x-axis provides an estimated melting temperature of K. This value is smaller than the recently estimated melting point of the TIP4P/2005 model of K Conde, Rovere, and Gallo (2017). For point charge models with similar dipole moments, the location of the melting point correlates with the quadrupole moment, Abascal and Vega (2007). As shown in Table 2, TIP4P/ has lower value of than the TIP4P/2005, explaining the shift in melting point by about K.
@@ IV.5 Transport properties
By virtue of the Stokes-Einstein relation, it is expected that the shear viscosity values should also be equivalently well reproduced by the TIP4P/ water model as it is inherently coupled to the self-diffusion coefficient. Indeed, it is the case as can be seen in Figure 6 where the results overlap with the data from Montero de Hijes et al. Montero de Hijes et al. (2018) for the TIP4P/2005 model. In Figure 6-b a magnified region is presented.
@@
In Figure 6-c, the relation of shear viscosity with pressure at three different temperatures is shown. Remarkably, the shear viscosities at K are around 10% too small as compared to the experimental data, however, the trend in which the viscosity reaches its minimum and increases again with pressure is preserved. Nevertheless, this is also the case for the TIP4P/2005 water model as demonstrated in Figure 2 in Ref. Montero de Hijes et al. (2018). As the temperature increases, the agreement with experimental results improves.
IV.6 Comparison of the models
@@
So far, for selected properties, we have seen that the TIP4P/ water model is essentially identical to the TIP4P/2005 force field in a wide range of properties, with the exception that the former also reproduces experimental values of the dielectric constant. Therefore, we are now in a position to perform a similar comparison as has been recently done by Sedano et al. (TIP4P/2005 vs OPC models), using the VA-test method Vega and Abascal (2011).
@@
The results for different properties are shown in Table 3. The comparison is done exactly as in Ref. Sedano, Blazquez, and Vega (2024), except for the heat capacities at constant pressure, which were not calculated in this work. All the results for the TIP4P/2005 and OPC water models are taken from that reference (except for the viscosity at 298 K which is taken from González and Abascal (2010)), whereas properties of the TIP4P/ are calculated by ourselves unless marked as an asterisk. In the latter case, the value for a given property is taken from Ref. Fuentes-Azcatl and Alejandre (2014). The tolerances allowed for variations of properties in the rating scheme are as those reported in Table 2 of Ref. Vega and Abascal (2011).
@@
Despite the wide range of properties tested in the VA benchmark, both models perform quite well. TIP4P/2005 is superior for the prediction of the melting and critical points, but is very similar to TIP4P/ for all other properties, including ice polymorph densities, liquid densities, shear viscosities and diffusion coefficients. The overall performance is almost the same for both models (7.59 for TIP4P/2005 vs 7.54 for TIP4P/) while the OPC is clearly inferior (with a score of 6.26). As a bonus, TIP4P/ predicts accurately the dielectric constants, compared to TIP4P/2005 which provides results that are ca. 25 % too small. Surprisingly, this large increase of is achieved with only a minor increase of the molecular dipole moment. The implication is that orientation correlations of the TIP4P/ model must be very different from those predicted by TIP4P/2005, yet the prediction of densities and transport properties in both models remain very similar. It would seem that changes in the Lennard-Jones parameters from one model to the other compensate somehow for the differences in charge distributions. The question then remains, whether one of the models is predicting the actual electric field about a water molecule better than the other, but this cannot be resolved from our results based on the properties of pure water.
@@
| Enthalpy of phase change (kJ/mol) | ||||
| 6.02 | 4.73 | 4.05 | 4.48 | |
| 44.02 | 50.17 | 53.97 | ||
@@
V Electrolyte properties of TIP4P/ model
In order to elucidate which of the two models is more consistent, it is required to test their performance beyond pure water properties. Particularly, a strategy to test which of the two charge distributions is electrostatically more significant is to test the models under strong electric fields, such as that generated locally in electrolyte solutions. A priori, one can tell that TIP4P/ could perform better for the solubilities of salts, as it predicts accurately the dielectric constant. However, if this agreement is at the cost of spoiling the force field, the remaining solution properties will not be predicted correctly.
V.1 Joung-Cheatham NaCl model
In the present work, we have decided to use the NaCl parameters of the force field that has been developed by Joung and Cheatham (JC) Joung and Cheatham (2008). The parameters used in conjunction with the TIP4P/ water model are shown in Table 4.
The motivation behind the choice of the JC model is three-fold. Firstly, the model has been parametrized to accurately reproduce the properties of the NaCl crystal and its melt. This constrains the parameters and avoids possible inconsistencies that could result from fitting to the solution properties alone. Secondly, as a corollary of the former comment, the JC model sets unit charges to the Na+ and Cl- ions, which is the textbook expectation for charge distributions of dissolved ions. Thirdly, the model parameters have not been optimized for either TIP4P/2005 or TIP4P/. This is convenient, as we can therefore check the performance of the water models without any a priori bias. The comparison will therefore allow us to check whether (i) the model electrolyte is transferable to other water models and (ii) reproducing the experimental values of dielectric constant improves the results.